# Restricted Maximum Likelihood Estimation in Generalized Linear Mixed Models

### v1

@@ 3.1 Linearization and Adjusted Dependent Variables
Wolfinger and O’Connell, (1993) noted their estimation procedure bears strong similarities to, and in some cases produce the same REML estimates as, the quasi-likelihood procedures of Breslow and Clayton, (1993) which we review in Section 3.3. Indeed, aside from whether $\phi$ is estimated versus assumed known, Breslow and Clayton, (1993) derived their estimation procedure using a Laplace-approximated marginal quasi-likelihood function, while Wolfinger and O’Connell, (1993) applied a Gaussian approximation directly to the adjusted dependent variables. Wolfinger and O’Connell, (1993) also noted that, in the special case of the GLMM studied in Schall, (1991), their method produces the same REML estimates of $\sigma^{2}_{k}$ as in Schall, (1991), with only minor differences in the fitting algorithm due to the use of the Newton-Raphson algorithm to maximize (8).
@@ 3.2 Linearization and the BLUP approach
McGilchrist, (1994) developed a class of approximate linearization methods based on the method of best linear unbiased predictions (BLUPs) used to fit general mixed-effects models, and exploited the connection between BLUPs and REML for LMMs, as developed by Henderson, (1963, 1975). Consider the model $g(\boldsymbol{\mu})=\boldsymbol{X}\boldsymbol{\beta}+\sum_{k=1}^{K}\boldsymbol{Z}_{k}\boldsymbol{u}_{k}$ and $\mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})=\mbox{diag}(\sigma^{2}_{1}\boldsymbol{A}_{1},\ldots,\sigma^{2}_{K}\boldsymbol{A}_{K})$. This is similar to the model in Schall, (1991), with the addition of known constant matrices $\boldsymbol{A}_{1},\ldots,\boldsymbol{A}_{K}$ that do not have to be diagonal in structure. From (3), the joint log-likelihood of the GLMM is
$\ell_{J}(\boldsymbol{\theta},\boldsymbol{\beta},\phi;\boldsymbol{y},\boldsymbol{u})=\log\{p(\boldsymbol{y}|\boldsymbol{u};\boldsymbol{\beta},\phi)\}-\frac{1}{2}\sum_{k=1}^{K}\left\{q_{k}\log(\sigma^{2}_{k})+\frac{1}{\sigma^{2}_{k}}\boldsymbol{u}^{\top}_{k}\boldsymbol{A}_{k}^{-1}\boldsymbol{u}_{k}\right\}\triangleq\ell_{1}+\ell_{2},$
where constants with respect to model parameters are omitted. Next, let $\boldsymbol{H}=-\nabla^{2}_{\boldsymbol{\beta},\boldsymbol{u}}\ell_{1}$ denote the observed information matrix for $\boldsymbol{\beta}$ and $\boldsymbol{u}$; the expected information matrix could also be used, although this is anticipated to make little difference in practice. Following arguments provided by McGilchrist and Aisbett, (1991) for LMMs, McGilchrist, (1994) then proposed replacing $\ell_{1}$ by a quadratic approximation around the estimators $(\tilde{\boldsymbol{\beta}}^{\top},\tilde{\boldsymbol{u}}^{\top})^{\top}=\arg\max_{\boldsymbol{\beta},\boldsymbol{u}}\ell_{1}$,
$\ell^{*}_{1}=-\frac{1}{2}\left[\begin{array}[]{c}\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\\[4.30554pt] \boldsymbol{u}-\tilde{\boldsymbol{u}}\end{array}\right]^{\top}\boldsymbol{H}\left[\begin{array}[]{c}\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\\[4.30554pt] \boldsymbol{u}-\tilde{\boldsymbol{u}}\end{array}\right]=-\frac{1}{2}(\boldsymbol{y}^{*}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u})^{\top}\boldsymbol{W}(\boldsymbol{y}^{*}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u}),$
@@
Viewing the approximate joint log-likelihood function $\ell^{*}_{J}(\boldsymbol{\theta},\boldsymbol{\beta},\phi;\boldsymbol{y},\boldsymbol{u})=\ell^{*}_{1}+\ell_{2}$ as one corresponding to that of an approximate linearized GLMM, McGilchrist, (1994) then applied a standard procedure for estimating LMMs as follows. Given $\boldsymbol{\theta}$ and $\phi$, we calculate the BLUP estimators of $\boldsymbol{\beta}$ and $\boldsymbol{u}$ by maximizing $\ell^{*}_{J}(\boldsymbol{\theta},\boldsymbol{\beta},\phi;\boldsymbol{y},\boldsymbol{u})$. This can be shown to be equivalent to solving the mixed model equations (7), with $\boldsymbol{\xi}^{\dagger}$ replaced by $\boldsymbol{y}^{*}$. Then, given $\boldsymbol{\beta}$ and $\boldsymbol{u}$, we compute updates for the REML estimates of $\sigma^{2}_{k}$ similar to those of Schall, (1991): for $k=1,\ldots,K$, we use the fixed-point update $\sigma^{2}_{k}\Leftarrow(\boldsymbol{u}_{k}^{\top}\boldsymbol{A}^{-1}_{k}\boldsymbol{u}_{k})/\{q_{k}-\sigma^{-2}_{k}\text{trace}(\boldsymbol{A}^{-1}_{k}\boldsymbol{T}_{kk})\}$. If $\phi$ requires estimation, then this can also be updated using a REML-like estimator.
Under the same choice of mean model, response distribution, and link function, the approximate linearization methods of McGilchrist, (1994) and Schall, (1991) are equivalent and produce the same REML estimates. This equivalence was noted by McGilchrist, (1994), who commented that the primary difference between the two methods lies in arguments by which the approximation linearization is obtained: Schall, (1991) and Wolfinger and O’Connell, (1993) considered linearization of the link function directly, while the linearization of McGilchrist, (1994) arises through approximating the first part of the joint log-likelihood function. In a small simulation study with binomial GLMMs, McGilchrist, (1994) showed their REML estimates exhibit less bias than a standard unrestricted maximum likelihood estimator, which they also derive, but may still present substantial bias if the true value of the variance component is large. On the other hand, McGilchrist, (1994) argued that the BLUP framework aims to target a broader class of mixed model applications. This is exemplified later by McGilchrist and Yau, (1995) who used the BLUP framework for REML estimation in a logistic GLMM with an AR(1) random component, Saei and McGilchrist, (1997) who used it in ordinal response mixed models, and Yau and McGilchrist, (1998) and Yau, (2001) who applied it to frailty models in survival analysis. The BLUP approach to calculating REML estimates has been further extended by Lee and Nelder, (1996) to hierarchical generalized linear models; see Section S.1 of the supplementary material for a more detailed review of this.
@@ 3.3 Linearization and Quasi-Likelihood Estimation
Quasi-likelihood methods exploit the relationship between the mean and variance of the responses to construct an objective function, without requiring the specification of a full distribution in the model (McCullagh and Nelder,, 1989, Chapter 9). In the setting of GLMMs, this was adopted by Breslow and Clayton, (1993) who proposed two quasi-likelihood methods, and subsequently developed REML estimation, for fitting GLMMs; see also Green, (1987) and Goldstein, (1991) for the same approach in the context of semi-parametric regression models and nonlinear multi-level mixed models, respectively.

### v2

@@ 3.1 Linearization and Adjusted Dependent Variables
Wolfinger and O’Connell (1993) noted their estimation procedure bears strong similarities to, and in some cases produces the same REML estimates as, the quasi-likelihood procedures of Breslow and Clayton (1993) which we review in Section 3.3. Indeed, aside from whether $\phi$ is estimated versus assumed known, Breslow and Clayton (1993) derived their estimation procedure using a Laplace-approximated marginal quasi-likelihood function, while Wolfinger and O’Connell (1993) applied a Gaussian approximation directly to the adjusted dependent variables. Wolfinger and O’Connell (1993) also noted that, in the special case of the GLMM studied in Schall (1991), their method produces the same REML estimates of $\sigma^{2}_{k}$ as in Schall (1991), with only minor differences in the fitting algorithm due to the use of the Newton-Raphson algorithm to maximize (8).
@@ 3.2 Linearization and the BLUP approach
McGilchrist (1994) developed a class of approximate linearization methods based on the method of best linear unbiased predictions (BLUPs) used to fit general mixed-effects models, and exploited the connection between BLUPs and REML for LMMs, as developed by Henderson (1963, 1975). Consider the model $g(\boldsymbol{\mu})=\boldsymbol{X}\boldsymbol{\beta}+\sum_{k=1}^{K}\boldsymbol{Z}_{k}\boldsymbol{u}_{k}$ and $\mbox{Cov}(\boldsymbol{u})=\boldsymbol{D}(\boldsymbol{\theta})=\mbox{diag}(\sigma^{2}_{1}\boldsymbol{A}_{1},\ldots,\sigma^{2}_{K}\boldsymbol{A}_{K})$. This is similar to the model in Schall (1991), with the addition of known constant matrices $\boldsymbol{A}_{1},\ldots,\boldsymbol{A}_{K}$ that do not have to be diagonal in structure. From (3), the joint log-likelihood of the GLMM is
$\ell_{J}(\boldsymbol{\beta},\phi,\boldsymbol{\theta};\boldsymbol{y},\boldsymbol{u})=\log\{p(\boldsymbol{y}|\boldsymbol{u};\boldsymbol{\beta},\phi)\}$
$\quad-\frac{1}{2}\sum_{k=1}^{K}\left\{q_{k}\log(\sigma^{2}_{k})+\frac{1}{\sigma^{2}_{k}}\boldsymbol{u}^{\top}_{k}\boldsymbol{A}_{k}^{-1}\boldsymbol{u}_{k}\right\}\triangleq\ell_{1}+\ell_{2},$
where constants with respect to model parameters are omitted. Next, let $\boldsymbol{H}=-\nabla^{2}_{\boldsymbol{\beta},\boldsymbol{u}}\ell_{1}$ denote the observed information matrix for $\boldsymbol{\beta}$ and $\boldsymbol{u}$; the expected information matrix could also be used, although this is anticipated to make little difference in practice. Following arguments provided by McGilchrist and Aisbett (1991) for LMMs, McGilchrist (1994) then proposed replacing $\ell_{1}$ by a quadratic approximation around the estimators $(\tilde{\boldsymbol{\beta}}^{\top},\tilde{\boldsymbol{u}}^{\top})^{\top}=\arg\max_{\boldsymbol{\beta},\boldsymbol{u}}\ell_{1}$,
$\ell^{*}_{1}$ $=-\frac{1}{2}\left[\begin{array}[]{c}\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\\[4.30554pt] \boldsymbol{u}-\tilde{\boldsymbol{u}}\end{array}\right]^{\top}\boldsymbol{H}\left[\begin{array}[]{c}\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\\[4.30554pt] \boldsymbol{u}-\tilde{\boldsymbol{u}}\end{array}\right]$
$=-\frac{1}{2}(\boldsymbol{y}^{*}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u})^{\top}\boldsymbol{W}(\boldsymbol{y}^{*}-\boldsymbol{X}\boldsymbol{\beta}-\boldsymbol{Z}\boldsymbol{u}),$
@@
Viewing the approximate joint log-likelihood function $\ell^{*}_{J}(\boldsymbol{\beta},\phi,\boldsymbol{\theta};\boldsymbol{y},\boldsymbol{u})=\ell^{*}_{1}+\ell_{2}$ as one corresponding to that of an approximate linearized GLMM, McGilchrist (1994) then applied a standard procedure for estimating LMMs as follows. Given $\boldsymbol{\theta}$ and $\phi$, we calculate the BLUP estimators of $\boldsymbol{\beta}$ and $\boldsymbol{u}$ by maximizing $\ell^{*}_{J}(\boldsymbol{\beta},\phi,\boldsymbol{\theta};\boldsymbol{y},\boldsymbol{u})$. This can be shown to be equivalent to solving the mixed model equations (7), with $\boldsymbol{\xi}^{\dagger}$ replaced by $\boldsymbol{y}^{*}$. Then, given $\boldsymbol{\beta}$ and $\boldsymbol{u}$, we compute updates for the REML estimates of $\sigma^{2}_{k}$ similar to those of Schall (1991): for $k=1,\ldots,K$, we use the fixed-point update $\sigma^{2}_{k}\Leftarrow(\boldsymbol{u}_{k}^{\top}\boldsymbol{A}^{-1}_{k}\boldsymbol{u}_{k})/\{q_{k}-\sigma^{-2}_{k}\text{trace}(\boldsymbol{A}^{-1}_{k}\boldsymbol{T}_{kk})\}$. If $\phi$ requires estimation, then this can also be updated using a REML-like estimator.
Under the same choice of mean model, response distribution, and link function, the approximate linearization methods of McGilchrist (1994) and Schall (1991) are equivalent and produce the same REML estimates. This equivalence was noted by McGilchrist (1994), who commented that the primary difference between the two methods lies in arguments by which the approximation linearization is obtained: Schall (1991) and Wolfinger and O’Connell (1993) considered linearization of the link function directly, while the linearization of McGilchrist (1994) arises through approximating the first part of the joint log-likelihood function. In a small simulation study with binomial GLMMs, McGilchrist (1994) showed their REML estimates exhibit less bias than a standard unrestricted maximum likelihood estimator, which they also derive, but may still present substantial bias if the true value of the variance component is large. On the other hand, McGilchrist (1994) argued that the BLUP framework aims to target a broader class of mixed model applications. This is exemplified later by McGilchrist and Yau (1995) who used the BLUP framework for REML estimation in a logistic GLMM with an AR(1) random component, Saei and McGilchrist (1997) who used it in ordinal response mixed models, and Yau and McGilchrist (1998) and Yau (2001) who applied it to frailty models in survival analysis. The BLUP approach to calculating REML estimates has been further extended by Lee and Nelder (1996) to HGLMs; see Section 7.
@@ 3.3 Linearization and Quasi-Likelihood Estimation
Quasi-likelihood methods exploit the relationship between the mean and variance of the responses to construct an objective function, without requiring the specification of a full distribution in the model (McCullagh and Nelder, 1989, Chapter 9). In the setting of GLMMs, this was adopted by Breslow and Clayton (1993) who proposed two quasi-likelihood methods, and subsequently developed REML estimation, for fitting GLMMs; see also Green (1987) and Goldstein (1991) for the same approach in the context of semi-parametric regression models and nonlinear multi-level mixed models, respectively.